# Scalable Computations for Generalized Mixed Effects Models with Crossed Random Effects Using Krylov Subspace Methods

### v1

@@ Abstract
Mixed effects models are widely used for modeling data with hierarchically grouped structures and high-cardinality categorical predictor variables. However, for high-dimensional crossed random effects, current standard computations relying on Cholesky decompositions can become prohibitively slow. In this work, we present novel Krylov subspace-based methods that address several existing computational bottlenecks. Among other things, we theoretically analyze and empirically evaluate various preconditioners for the conjugate gradient and stochastic Lanczos quadrature methods, derive new convergence results, and develop computationally efficient methods for calculating predictive variances. Extensive experiments using simulated and real-world data sets show that our proposed methods scale much better than Cholesky-based computations, for instance, achieving a runtime reduction of approximately two orders of magnitudes for both estimation and prediction. Moreover, our software implementation is up to 10’000 times faster and more stable than state-of-the-art implementations such as lme4 and glmmTMB when using default settings. Our methods are implemented in the free C++ software library GPBoost with high-level Python and R packages.
@@ 1 Introduction
Mixed effects models are widely used in various scientific disciplines for modeling data with hierarchically grouped structures (Laird and Ware, 1982; Pinheiro and Bates, 2000). High-dimensional crossed random effects occur frequently in practice, e.g., when modeling ratings in recommender systems with categorical grouping variables corresponding to customers and products (Gao and Owen, 2017; Ghosh et al., 2022a; Simchoni and Rosset, 2023) or, in general, when there are high-cardinality categorical predictor variables (Simchoni and Rosset, 2021; Sigrist, 2023b). However, current standard Cholesky decomposition-based computations can become prohibitively slow for high-dimensional crossed random effects. For instance, the cost of computing Cholesky factors of sparse matrices can be of cubic complexity in the dimension of the random effects (Pandolfi et al., 2024).
@@
Various methods have been proposed to overcome this computational bottleneck, including method of moments estimation (Gao and Owen, 2017, 2020), collapsed Gibbs samplers (Papaspiliopoulos et al., 2020, 2023), backfitting algorithms (Ghosh et al., 2022b, a), composite likelihoods (Bellio et al., 2023; Xu et al., 2023), and variational inference (Xu et al., 2023; Goplerud et al., 2025). Furthermore, the preconditioned conjugate gradient (CG) method (Saad, 2003) with relatively simple diagonal or block-diagonal preconditioners has been applied to solve high-dimensional mixed model equations (Strandén and Lidauer, 1999, 2001; Tsuruta et al., 2001; Taskinen et al., 2017; Garrick et al., 2019; Vandenplas et al., 2018). Recently, Pandolfi et al. (2024) have used the CG method in Gibbs sampling for Bayesian generalized linear mixed models (GLMMs). For Gaussian likelihoods and the special case where all random effects have the same variance, Border and Becker (2019) and Cheng et al. (2023) use stochastic Lanczos quadrature (SLQ) (Ubaru et al., 2017) without preconditioning in combination with a shift-invariance property to calculate log-determinants in log-likelihoods. However, without preconditioning, SLQ-approximated log-determinants have high variance and are thus inaccurate; see our theoretical and empirical analyses below. Overall, there are two major unsolved computational challenges: (i) efficiently computing log-determinants in log-likelihoods and their gradients, and (ii) solving high-dimensional sparse linear systems with many right-hand sides for prediction.
In this article, we show how Krylov subspace methods, such as the preconditioned CG and SLQ methods, can be used for computationally efficient parameter estimation and the calculation of posterior predictive distributions for generalized mixed effects models (GMMs), a term we use to denote mixed effects models with linear and non-linear fixed effects functions and general response variable distributions. For non-Gaussian likelihoods, marginal likelihoods are approximated using the Laplace approximation since it is computationally efficient and converges asymptotically to the correct quantity. Supporting this argument, we show in Figure 1 the estimated variance parameter of one random effect in a model with two crossed random effects obtained with a Laplace approximation for varying numbers of repeated observations $d$ per random effect realization on simulated data with a Bernoulli likelihood and the setting described in Section 5.1 with $m=4^{\prime}000$. Estimation is done using the Krylov subspace methods introduced in this paper and repeated on $100$ simulated data sets for every $d$. Figure 1 shows that the variance parameter estimator has a small downward bias which vanishes as $d$ grows.
Figure 1: Estimated variance parameter $\sigma^{2}_{1}$ obtained with a Laplace approximation vs. different numbers of repeated observations $d$ per random effect realization for binary data. The red rhombi represent means and the whiskers are ±2$\pm$ 2$\times$ 2 × standard errors. The dashed line indicates the true parameter $\sigma^{2}_{1}=0.25$.
Krylov subspace methods allow for fast computations as they rely on matrix-vector multiplications with sparse matrices which can be trivially parallelized. We use the preconditioned CG and SLQ methods to solve sparse linear systems and to calculate log-determinants in log-likelihoods, respectively. Gradients of log-marginal likelihoods are calculated using stochastic trace estimation (STE) with almost no computational overhead once likelihoods are evaluated. We theoretically and empirically analyze several preconditioners with regard to the accuracy of SLQ approximations and the CG method’s convergence speed. Our theoretical results show that approximations with the SLQ method are expected to be accurate and that the CG method converges rapidly. In addition, we provide theoretical and empirical results showing that the symmetric successive over-relaxation (SSOR) preconditioner is superior compared to the popular diagonal preconditioner for both the CG and the SLQ methods. Furthermore, we introduce efficient and accurate simulation-based methods for calculating predictive variances when the number of prediction points is large. In experiments with simulated and real-world data, we obtain a reduction in runtime of approximately two orders of magnitude compared to Cholesky-based computations for estimation and prediction while having essentially identical accuracy. Moreover, our software implementation in the GPBoost library is up to 10’000 times faster and more stable than the state-of-the-art lme4 and glmmTMB packages when using default settings.
@@ 2 Preliminaries on Generalized Mixed Effects Models
$\mu=F(X)+Zb,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ b\sim N(0,\Sigma),$
where $X\in\mathbb{R}^{n\times p}$ contains predictor variables. In generalized linear mixed effects models (GLMMs), the fixed effects function is linear, $F(X)=X\beta$, but $F(\cdot)$ can also be modeled using machine learning methods such as regression trees (Hajjem et al., 2011; Sela and Simonoff, 2012; Fu and Simonoff, 2015), random forests (Hajjem et al., 2014), tree-boosting (Sigrist, 2022, 2023a), or deep neural networks (Simchoni and Rosset, 2021, 2023; Avanzi et al., 2024). The latent Gaussian variables $b=(b_{1}^{T},b_{2}^{T},\dots,b_{K}^{T})^{T}\in\mathbb{R}^{m}$ consist of $K$ grouped random effects $b_{j}\in\mathbb{R}^{m_{j}}$, $j\in\{1,\dots,K\}$, $m=\sum_{j=1}^{K}m_{j}$, and the covariance matrix $\Sigma\in\mathbb{R}^{m\times m}$ is diagonal and depends on a set of variance parameters $\theta=\{\sigma_{1}^{2},\sigma_{2}^{2},\dots,\sigma_{K}^{2}\}$:
@@
In addition, $Z=(Z_{1},Z_{2},\dots,Z_{K})\in\mathbb{R}^{n\times m}$, $Z_{j}\in\mathbb{R}^{n\times m_{j}}$, is usually a binary incidence matrix that maps the random effects to the corresponding observations, but it can also contain predictor variables when modeling random coefficients. We denote such a model as a generalized mixed effects model (GMM). Since we use the Laplace approximation, we further assume that $p(y|\mu,\xi)$ is log-concave in $\mu$.
@@ 2.1 Parameter estimation
[…]
$L^{LA}(y|F,\theta,\xi)=-\log p(y|\mu^{*},\xi)+\frac{1}{2}{b^{*}}^{T}\Sigma^{-1}b^{*}+\frac{1}{2}\log\det\left(\Sigma\right)+\frac{1}{2}\log\det\left(\Sigma^{-1}+Z^{T}WZ\right),$ (1)
where $\mu^{*}=F(X)+Zb^{*}$, $b^{*}=\operatorname*{argmax}_{b}\log p(y|\mu,\xi)-\frac{1}{2}b^{T}\Sigma^{-1}b$ is the mode of $p(y|\mu,\xi)p(b|\theta)$, and $W\in\mathbb{R}^{n\times n}$ is diagonal with $W_{ii}=-\frac{\partial^{2}\log p(y_{i}|\mu_{i},\xi)}{\partial\mu_{i}^{2}}\Big{|}_{\mu=\mu^{*}}$. The mode $b^{*}$ is typically found with Newton’s method (Williams and Rasmussen, 2006), and one iteration is given by
$b^{*t+1}=b^{*t}+\left(Z^{T}WZ+\Sigma^{-1}\right)^{-1}\left(Z^{T}\frac{\partial\log p(y|\mu^{*t},\xi)}{\partial b}-\Sigma^{-1}b^{*t}\right),\leavevmode\nobreak\ \leavevmode\nobreak\ t=0,1,\dots$ (2)
If a first- or second-order optimization method is used for minimizing $L^{LA}(y|F,\theta,\xi)$, gradients with respect to $\theta$, $F$, and $\xi$ are needed. These gradients can be found, e.g., in Sigrist (2023a). Note that gradients with respect to $F$ are used, for instance, for GLMMs since $\frac{\partial L^{LA}(y|F,\theta,\xi)}{\partial\beta}=X^{T}\frac{\partial L^{LA}(y|F,\theta,\xi)}{\partial F}$ when $F(X)=X\beta$.
@@
$y=F(X)+Zb+\epsilon,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \epsilon\sim N(0,W^{-1}),\leavevmode\nobreak\ \leavevmode\nobreak\ W=\frac{1}{\sigma^{2}}I_{n},$
where $\sigma^{2}$ is the error variance and $I_{n}\in\mathbb{R}^{n\times n}$ an identity matrix. In this case, the Laplace approximation is exact, and the negative log-marginal likelihood can be equivalently calculated without determining the mode $b^{*}$ as follows:
$L(y|F,\theta,\xi)=\frac{n}{2}\log(2\pi)+\frac{1}{2}\log\det(\Psi)+\frac{1}{2}(y-F(X))^{T}\Psi^{-1}(y-F(X)),\leavevmode\nobreak\ \leavevmode\nobreak\ \Psi=Z\Sigma Z^{T}+W^{-1}.$ (3)
Since the dimension of the random effects $m$ is usually smaller than the number of observations $n$, one uses the Woodbury identity
@@
$\frac{\partial L(y|F,\theta,\xi)}{\partial\theta_{k}}=\frac{1}{2}\operatorname*{tr}\left((\Sigma^{-1}+Z^{T}WZ)^{-1}\frac{\partial\Sigma^{-1}}{\partial\theta_{k}}+\Sigma^{-1}\frac{\partial\Sigma}{\partial\theta_{k}}\right)-\frac{1}{2}(y-F(X))^{T}\Psi^{-1}\frac{\partial\Psi}{\partial\theta_{k}}\Psi^{-1}(y-F(X)).$ (6)
@@ 2.2 Prediction
Consider the goal of making predictions for the latent variables
@@
where $X_{p}\in\mathbb{R}^{n_{p}\times p}$ is the predictor variable matrix of the predictions, $b_{p}\in\mathbb{R}^{m_{p}}$ are new random effects which are not in the training data random effects $b\in\mathbb{R}^{m}$, the matrix $Z_{po}\in\mathbb{R}^{n_{p}\times m}$ relates $b$ to $\mu_{p}$, and the matrix $Z_{pp}\in\mathbb{R}^{n_{p}\times m_{p}}$ relates the new random effects $b_{p}$ to $\mu_{p}$. We have
$\begin{split}\begin{pmatrix}b\\ \mu_{p}\end{pmatrix}&=\begin{pmatrix}0\\ F(X_{p})\end{pmatrix}+\begin{pmatrix}(I_{m},0_{m\times m_{p}})\\ (Z_{po},Z_{pp})\end{pmatrix}\begin{pmatrix}b\\ b_{p}\end{pmatrix}\sim\mathcal{N}\left(\begin{pmatrix}0\\ F(X_{p})\end{pmatrix},\begin{pmatrix}\Sigma&\Sigma Z_{po}^{T}\\ Z_{po}\Sigma&Z_{po}\Sigma Z_{po}^{T}+Z_{pp}\Sigma_{p}Z_{pp}^{T}\end{pmatrix}\right),\end{split}$ (9)
where $(I_{m},0_{m\times m_{p}})\in\mathbb{R}^{m\times(m+m_{p})}$, $I_{m}\in\mathbb{R}^{m\times m}$ is an identity matrix, $0_{m\times m_{p}}\in\mathbb{R}^{m\times m_{p}}$ is a matrix of zeros, and $\Sigma_{p}=\text{Cov}(b_{p})$. Predictions for $\mu_{p}$ are obtained using the posterior predictive distribution $p(\mu_{p}|y,\theta,\xi)$ which is given by
@@
This can be derived using the fact that a Laplace approximation for $p(y|F,\theta,\xi)$ is equivalent to the approximation $p(b|y,\theta,\xi)\approx\mathcal{N}\left(b^{*},\left(\Sigma^{-1}+Z^{T}WZ\right)^{-1}\right)$ since $p(y|F,\theta,\xi)=p(y|F,b,\xi)p(b|\theta)/p(b|y,\theta,\xi)$, and by applying standard results for conditional distributions of multivariate Gaussian distributions.
For the posterior predictive distribution $p(y_{p}|y,\theta,\xi)$ of the observable response variables $y_{p}$, $y_{p}|\mu_{p}\sim p(y_{p}|\mu_{p},\xi)$, the integral
[…]
must be additionally calculated. This is analytically intractable for most likelihoods but can be approximated using numerical integration or by simulating from $p(\mu_{p}|y,\theta,\xi)\approx\mathcal{N}\left(\omega_{p},\Omega_{p}\right)$.
@@
and $p(y_{p}|y,\theta,\xi)$ is obtained from $p(\mu_{p}|y,\theta,\xi)$ by adding a diagonal matrix with the error variance to $\Omega_{p}$.
Note that, while the above predictive distributions might be less frequently used in classical statistical applications of GLMMs, they are important, e.g., for mixed effects machine learning models (Sigrist, 2022; Simchoni and Rosset, 2023) for high-cardinality categorical variables when probabilistic predictions are required.

### v2

# Scalable Krylov Subspace Methods for Generalized Mixed Effects Models with Crossed Random Effects

@@ Abstract
Mixed-effects models are widely used to model data with hierarchical grouping structures and high-cardinality categorical predictor variables. However, for high-dimensional crossed random effects, sparse Cholesky decompositions, the current standard approach, can become prohibitively slow. In this work, we present Krylov subspace-based methods that address these computational bottlenecks and analyze them both theoretically and empirically. In particular, we derive new results on the convergence and accuracy of the preconditioned stochastic Lanczos quadrature and conjugate gradient methods for mixed-effects models, and we develop scalable methods for calculating predictive variances. In experiments with simulated and real-world data, the proposed methods yield speedups by factors of up to about 10,000 and are numerically more stable than Cholesky-based computations as implemented in state-of-the-art packages such as lme4 and glmmTMB. Our methodology is available in the open-source C++ software library GPBoost, with accompanying high-level Python and R packages.
@@ 1 Introduction
Mixed effects models are widely used in various scientific disciplines to analyze data with hierarchically grouped structures (pinheiro2000linear; mcculloch2004generalized). High-dimensional crossed random effects occur frequently in practice, for instance, when modeling ratings in recommender systems with categorical grouping variables corresponding to customers and products (gao2017efficient; ghosh2022backfitting; simchoni2023integrating) or, in general, when there are high-cardinality categorical predictor variables (simchoni2021using; sigrist2023comparison). However, current standard sparse Cholesky decomposition-based computations can become prohibitively slow for high-dimensional crossed random effects. For instance, the cost of computing Cholesky factors has up to cubic complexity in the dimension of the random effects (pandolfi2024conjugate).
@@
Various methods have been proposed to overcome this computational bottleneck, including method of moments estimation (gao2017efficient; gao2020estimation), collapsed Gibbs samplers (papaspiliopoulos2020scalable; papaspiliopoulos2023scalable), backfitting algorithms (ghosh2022scalable; ghosh2022backfitting), composite likelihoods (bellio2023consistent; xu2023gaussian), and variational inference (xu2023gaussian; goplerud2024partially). Furthermore, the preconditioned conjugate gradient (CG) method (saad2003iterative) with relatively simple diagonal or block-diagonal preconditioners has been applied to solve high-dimensional mixed model equations (stranden1999solving; stranden2001parallel; tsuruta2001use; taskinen2017single; garrick2019alternative; vandenplas2018deflated). Recently, pandolfi2024conjugate have analyzed the CG method with diagonal preconditioning for Bayesian generalized linear mixed models (GLMMs). For Gaussian likelihoods and the special case where all random effects have the same variance, border2019stochastic and cheng2023slemm use stochastic Lanczos quadrature (SLQ) (ubaru2017fast) without preconditioning in combination with a shift-invariance property to calculate log-determinants in log-marginal likelihoods. However, without preconditioning, SLQ-approximated log-marginal likelihoods have high variances and are thus inaccurate; see our theoretical and empirical analyses below. An accurate and scalable method for calculating log-determinants with a theoretical understanding of its convergence properties is thus currently missing. Furthermore, to our knowledge, there exists no scalable method for calculating predictive variances when the number of prediction points is large.
In this article, we introduce Krylov subspace methods for computationally efficient parameter estimation and the calculation of posterior predictive distributions for generalized mixed effects models with Gaussian and non-Gaussian likelihoods. Krylov subspace methods allow for fast computations as they rely on matrix-vector multiplications with sparse matrices, which can be trivially parallelized. We use the preconditioned CG and SLQ methods to solve sparse linear systems and to calculate log-determinants in log-likelihoods, respectively. Gradients of log-marginal likelihoods are calculated using stochastic trace estimation (STE), with almost no computational overhead once marginal likelihoods are computed.
We theoretically and empirically analyze the proposed methods. In particular, we derive novel results on extremal eigenvalues, condition numbers, and thus the convergence behavior of the SLQ method for the symmetric successive over-relaxation (SSOR) and diagonal preconditioners, as well as for the unpreconditioned case. Moreover, we present new results on the effective condition numbers and, hence, on the convergence of the CG method for the SSOR preconditioner, extending the results of pandolfi2024conjugate for the diagonal preconditioner. Our theoretical and empirical results show that the SSOR preconditioner yields more accurate SLQ-approximated log-marginal likelihoods and faster convergence of the CG method compared to the diagonal preconditioner and the unpreconditioned case. Furthermore, we introduce scalable and accurate simulation-based methods for calculating predictive variances when the number of prediction points is large. In experiments with simulated and real-world data, the proposed methods yield speedups by factors of up to about 10,000 and are numerically more stable than Cholesky-based computations as implemented in state-of-the-art packages such as lme4 (bates2015fitting) and glmmTMB (glmmtmb_cite), while maintaining essentially identical accuracy.
@@ 2 Preliminaries on Generalized Mixed Effects Models
$\mu=F(X)+Zb,\penalty 10000\ \penalty 10000\ \penalty 10000\ b\sim N(0,\Sigma),$
where $X\in\mathbb{R}^{n\times p}$ contains predictor variables. In generalized linear mixed effects models (GLMMs), the fixed effects function is linear, $F(X)=X\beta$, but $F(\cdot)$ can also be modeled using machine learning methods such as regression trees (hajjem2011mixed; sela2012re; fu2015unbiased), random forests (hajjem2014mixed), tree-boosting (sigrist2022gaussian; sigrist2022latent), or deep neural networks (simchoni2021using; simchoni2023integrating; avanzi2024machine). For notational simplicity, we will denote the parameters of this function $F(\cdot)$ by $\beta$, regardless of whether $F(\cdot)$ is linear or not. The latent Gaussian variables $b=(b_{1}^{T},b_{2}^{T},\dots,b_{K}^{T})^{T}\in\mathbb{R}^{m}$ consist of $K$ grouped random effect components $b_{k}\in\mathbb{R}^{m_{k}}$, $k\in\{1,\dots,K\}$, $m=\sum_{k=1}^{K}m_{k}$, and the covariance matrix $\Sigma\in\mathbb{R}^{m\times m}$ is diagonal and depends on a set of variance parameters $\theta=\{\sigma_{1}^{2},\sigma_{2}^{2},\dots,\sigma_{K}^{2}\}$:
@@
In addition, $Z=(Z_{1},Z_{2},\dots,Z_{K})\in\mathbb{R}^{n\times m}$, $Z_{j}\in\mathbb{R}^{n\times m_{k}}$, is usually a binary incidence matrix that maps the random effects $b$ to the corresponding observations, but it can also contain predictor variables when modeling random coefficients. We denote such a model as a generalized mixed effects model. Note that the diagonal assumption for $\Sigma$ is not crucial for the methods presented in the following and could be easily relaxed, e.g., to account for prior correlation among random slopes and intercepts.
@@ 2.1 Parameter estimation
For parameter estimation, the marginal likelihood $p(y|\beta,\theta,\xi)=\int p(y|\mu,\xi)p(b|\theta)db$ is typically maximized. For non-Gaussian likelihoods, there is no analytic expression for this marginal likelihood and an approximation has to be used. We use the Laplace approximation since it is computationally efficient and widely used in popular GLMM software libraries such as lme4 (bates2015fitting) and glmmTMB (glmmtmb_cite). For the Laplace approximation, we further assume that the likelihood $p(y|\mu,\xi)$ is log-concave and two times continuously differentiable in $\mu$. A Laplace approximation to the negative log-marginal likelihood $\log(p(y|\beta,\theta,\xi))$ modulo constant terms is given by
$L^{LA}(y|\beta,\theta,\xi)=-\log p(y|\mu^{*},\xi)+\frac{1}{2}{b^{*}}^{T}\Sigma^{-1}b^{*}+\frac{1}{2}\log\det\left(\Sigma\right)+\frac{1}{2}\log\det\left(\Sigma^{-1}+Z^{T}WZ\right),$ (1)
where $\mu^{*}=F(X)+Zb^{*}$, $b^{*}=\operatorname*{argmax}_{b}\log p(y|\mu,\xi)-\frac{1}{2}b^{T}\Sigma^{-1}b$ is the mode of $p(y|\mu,\xi)p(b|\theta)$, and $W\in\mathbb{R}^{n\times n}$ is diagonal with $W_{ii}=-\frac{\partial^{2}\log p(y_{i}|\mu_{i},\xi)}{\partial\mu_{i}^{2}}\Big|_{\mu=\mu^{*}}$. The mode $b^{*}$ is typically found with Newton’s method (williams2006gaussian), and one iteration is given by
$b^{*t+1}=b^{*t}+\left(Z^{T}WZ+\Sigma^{-1}\right)^{-1}\left(Z^{T}\frac{\partial\log p(y|\mu^{*t},\xi)}{\partial b}-\Sigma^{-1}b^{*t}\right),\penalty 10000\ \penalty 10000\ t=0,1,\dots$ (2)
If a first- or second-order optimization method is used for minimizing $L^{LA}(y|\beta,\theta,\xi)$, gradients with respect to $\theta$, $F=F(X)\in\mathbb{R}^{n}$, and $\xi$ are needed. These gradients can be found, e.g., in sigrist2022latent. Note that gradients with respect to $F\in\mathbb{R}^{n}$ are used, for instance, for GLMMs since $\frac{\partial L^{LA}(y|\beta,\theta,\xi)}{\partial\beta}=X^{T}\frac{\partial L^{LA}(y|\beta,\theta,\xi)}{\partial F}$ by the chain rule when $F(X)=X\beta$.
@@
$y=F(X)+Zb+\epsilon,\penalty 10000\ \penalty 10000\ \penalty 10000\ \epsilon\sim N(0,W^{-1}),\penalty 10000\ \penalty 10000\ W=\frac{1}{\sigma^{2}}I_{n},$
where $\sigma^{2}$ is the error variance and $I_{n}\in\mathbb{R}^{n\times n}$ an identity matrix. In this case, the Laplace approximation is exact, and the negative log-marginal likelihood is given by
$L(y|\beta,\theta,\xi)=\frac{n}{2}\log(2\pi)+\frac{1}{2}\log\det(\Psi)+\frac{1}{2}(y-F(X))^{T}\Psi^{-1}(y-F(X)),\penalty 10000\ \penalty 10000\ \Psi=Z\Sigma Z^{T}+W^{-1}.$ (3)
Since the dimension of the random effects $m$ is usually smaller than the number of observations $n$ and for sparsity reasons, one uses the Woodbury identity
@@
$\frac{\partial L(y|\beta,\theta,\xi)}{\partial\theta_{k}}=\frac{1}{2}\operatorname*{tr}\left((\Sigma^{-1}+Z^{T}WZ)^{-1}\frac{\partial\Sigma^{-1}}{\partial\theta_{k}}+\Sigma^{-1}\frac{\partial\Sigma}{\partial\theta_{k}}\right)-\frac{1}{2}(y-F(X))^{T}\Psi^{-1}\frac{\partial\Psi}{\partial\theta_{k}}\Psi^{-1}(y-F(X)).$ (6)
@@ 2.2 Prediction
Predictions for latent variables $\mu_{p}\in\mathbb{R}^{n_{p}}$ and response variables $y_{p}\in\mathbb{R}^{n_{p}}$ are made using the posterior predictive distributions $p(\mu_{p}|y)$ and $p(y_{p}|y)$, respectively. Note that predictions can be made for random effects $b$ corresponding to levels present in the training data and new random effects $b_{p}\in\mathbb{R}^{m_{p}}$ that did not occur in the observed data. We denote by $Z_{po}\in\mathbb{R}^{n_{p}\times m}$ the matrix that relates the random effects $b$ from the training data to the prediction points, and by $Z_{pp}\in\mathbb{R}^{n_{p}\times m_{p}}$ the matrix that relates the new random effects $b_{p}$ to the prediction points. Using this notation, we have
@@
where $X_{p}\in\mathbb{R}^{n_{p}\times p}$ is the predictor variable matrix of the prediction points, and
@@
where $(I_{m},0_{m\times m_{p}})\in\mathbb{R}^{m\times(m+m_{p})}$, $I_{m}\in\mathbb{R}^{m\times m}$ is an identity matrix, $0_{m\times m_{p}}\in\mathbb{R}^{m\times m_{p}}$ is a matrix of zeros, and $\Sigma_{p}=\text{Cov}(b_{p})$. While predictive distributions are less commonly used in classical statistical applications of GLMMs, they are important, for example, in mixed-effects machine learning models (sigrist2022gaussian; simchoni2023integrating), where probabilistic predictions are required.
For the latent variable $\mu_{p}$, the posterior predictive distribution $p(\mu_{p}|y)$ is given by
$p(\mu_{p}|y)=\int p(\mu_{p}|b)p(b|y)db\approx\mathcal{N}\left(\omega_{p},\Omega_{p}\right),$
@@
This can be derived using the fact that a Laplace approximation for $p(y|\beta,\theta,\xi)$ is equivalent to the approximation $p(b|y)\approx\mathcal{N}\left(b^{*},\left(\Sigma^{-1}+Z^{T}WZ\right)^{-1}\right)$ since $p(b|y)=p(y|b)p(b)/p(y)$, and by applying standard results for conditional distributions of multivariate Gaussian distributions.
For predicting the observable response variables $y_{p}$, $y_{p}|\mu_{p}\sim p(y_{p}|\mu_{p})$, an additional integral must be calculated:
$p(y_{p}|y)=\int p(y_{p}|\mu_{p})p(\mu_{p}|y)d\mu_{p}.$ (13)
This is analytically intractable for most likelihoods but can be approximated using numerical integration or by simulating from $p(\mu_{p}|y)\approx\mathcal{N}\left(\omega_{p},\Omega_{p}\right)$. Note that, for notational simplicity, we omit the dependence of $p(\mu_{p}|y)$, $p(b|y)$, $p(\mu_{p}|b)$, $p(y_{p}|y)$, and $p(y_{p}|\mu_{p})$ on $\beta$, $\theta$, and $\xi$.
@@
and $p(y_{p}|y)$ is obtained from $p(\mu_{p}|y)$ by adding the error variance to the diagonal of $\Omega_{p}$.